a rectangular plot of land is 50 yards shorter than 4 times the width. find the length if the perimeter is 900 yards
Answer by waynest(281)

length (L) = 4w - 50
width = w
perimeter = 900
2L + 2w = p
2(4w - 50) + 2w = 900
8w - 100 + 2w = 900
10w - 100 + 100 = 900 + 100
10w = 1000
10w/10 = 1000/10
w = 100
width = 100 yards
lenth = 4(100) - 50
= 400 - 50
= 350 yards
check:
2(100) + 2(350) = 900
200 + 700 = 900
900 = 900